# False vacuum decay in triamond lattice gauge theory

### v1

@@ I Introduction
The classic picture of false vacuum decay is a double-well potential energy function where the shallower well (called the false vacuum) is separated from the deeper well (called the true vacuum) by an energy barrier [1, 2]. A system that begins in the false vacuum can later be found in the true vacuum with some probability. For the corresponding scenario in a quantum field theory, the decay rate can be calculated by using a semi-classical approximation and imaginary time, but that approach does not explain the physical mechanism of the decay.
In contrast, real-time simulations would allow direct observation of the vacuum decay mechanism, perhaps showing bubbles of true vacuum appearing and nucleating to displace the false vacuum. Various approaches are being explored by several authors [3, 4, 5, 6, 7, 8]. In the present work, we take a first step toward using lattice gauge theory for the real-time simulation of false vacuum decay. We use a pure gauge theory; no matter fields are required.
@@
A promising approach to real-time physics is to rewrite the theory as a Hamiltonian problem, keeping time as a continuous variable and discretizing only the spatial components onto a three-dimensional lattice [11]. The length of the state vector grows exponentially with the lattice size but, in principle, it can be stored efficiently in the qubit register of a quantum computer [12]. Unfortunately, the non-Abelian version of Gauss’s law must now be imposed as a constraint that is external to the Hamiltonian, and finding the best approach to this issue is a high-priority topic for several research groups [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Nevertheless, non-Abelian theories have already been put successfully onto quantum computers [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], though the lattices are necessarily small and typically have just one or two spatial dimensions.
@@
In the absence of matter fields, gauge invariance requires any flow of non-Abelian charges (such as the colors of QCD) to form closed paths on the lattice. Localized particles called glueballs are gauge-invariant objects described by a superposition of localized paths. If the lattice has periodic boundary conditions, there can also be paths that go all the way around the lattice to close on themselves from behind as shown in Fig. 1, and these are called torelons [49, 50, 51, 52, 53, 54, 55, 56]. A torelon is a closed path of color flux that is not homotopic to a point. Torelons become non-dynamical heavy states as the lattice volume is taken to infinity, but on small lattices they are central to the physics.
Figure 1: A closed path of round-the-world color flux, called a torelon, in a cubic spatial volume with periodic boundary conditions. (a) One horizontal torelon. (b) One vertical torelon. (c) One diagonal torelon.
@@ II The triamond lattice
A primary motivation for choosing a triamond lattice rather than a cubic lattice is to reduce the number of gauge links touching each site from six to three [38]. Having only three gauge links at a site means the combined SU(2) value of any pair of links is determined uniquely by the value of the third gauge link. In contrast, the combined SU(2) value of two links on a cubic lattice is not unique because the other four links can be combined in many ways, and this means extra qubits would be required at each lattice site to fully specify the quantum state.
@@
Figure 2: (a) A triamond lattice comprising three unit cells with periodic boundary conditions in all three directions. Twelve of the 36 gauge links are numbered so they can be referenced in the text. (b) A flattened map of the same triamond lattice with the same 12 labels. (c) The same flattened map with two plaquettes shown in thick lines, namely a 10-sided plaquette at 1,2,3,4 and an 8-sided plaquette at 9,11. (d) Thick black paths show a horizontal, diagonal and vertical torelon in the left, center and right unit cells respectively.
@@
On our small triamond lattice, some of the elementary plaquettes wrap around a periodic boundary and affect the same link twice. In the special case where the twice-affected link retains its original value, the result is pair creation of two torelons that we will refer to as an 8-sided plaquette. Fig. 2(c) shows examples of 10-sided and 8-sided plaquettes.
@@
The mapping of a triamond lattice to a ladder of square paths is quite convenient. The three-dimensional aspects of the triamond lattice appear through the 10-sided and 8-sided plaquettes that involve more than just nearest-neighbour interactions among the squares.
@@ III The torelon spectrum
Figure 3: The smallest eigenvalues of SU(2) gauge theory on a triamond lattice having three unit cells in a row and periodic boundary conditions. The gauge theory is truncated such that each gauge link has $j\in\{0,\frac{1}{2}\}$ and the gauge coupling is $g=1$. Numerical values from this graph are provided in App. A.
@@
The 12 lowest energy states in the diagonal torelon sector tell an important story about how momentum is handled on a triamond lattice. These 12 lowest states are the energy eigenstates that correspond to placing a diagonal torelon at any of the 12 allowed positions on the lattice. In the ladder map of Fig. 2(b), they appear as superpositions of a single six-link rectangle at each of the 12 possible locations. Notice that half of these rectangles are doubly-blue and non-yellow while the other half are non-blue and doubly-yellow.
The translation along the 12-step ladder of Fig. 2(b) involves a $\pi/2$ rotation with each step along the $\hat{j}$ direction and is called a screw-translation. The Fourier transform of this screw-translation results in eigenstates of screw-momentum that are also eigenstates of energy, and these are precisely the states shown in Fig. 3. Importantly, the 12 values of screw-momentum contain six values of linear momentum for the non-yellow torelons and six values for the non-blue torelons. On the original triamond lattice, the non-blue torelon lies along a lattice diagonal that is spatially orthogonal to the non-yellow torelon. A non-blue torelon path is sketched in Fig. 2(d).
@@ IV Imaginary time evolution
Although the triamond lattice for three unit cells can be mapped onto the ladder of Fig. 2(b), the three-dimensional nature of the triamond structure remains evident through the large plaquettes shown in Fig. 2(c). A simpler system is obtained by using the ladder directly as a quasi-one-dimensional lattice having only square plaquettes and no connection to the three-dimensional triamond structure. This ladder of simple square plaquettes has been used in several quantum computations for SU($N$) gauge theory [25, 26, 27, 28, 29]. Imaginary time evolution [60] for a non-periodic ladder of only three rungs was run on a quantum computer recently [38]. Here we will study the 12-rung periodic ladder.
@@
For the 12-qubit Hamiltonian of Eq. (9), the most general expression has many terms but we might expect only a subset of them to dominate the physics. In particular, since $H_{\square}$ has only single-qubit terms, adjacent-pair terms, and adjacent-triple terms, we anticipate that $A$ will be predominantly local. Our ansatz has only single-qubit and adjacent-pair terms,
@@
Notice that each qubit $j$ gets its own set of coefficients $(a_{\bullet})_{j}$, thus allowing our error mitigation to make no assumption about individual physical qubits having similar noise profiles. The coefficients $(a_{\bullet})_{j}$ are determined from state tomography as explained in detail in App. B. The exponentiation of each term in $A$ is accomplished by using
@@
and their numerical values are different at each time step in the quantum circuit. Values from previous time steps are stored in a list for reuse, and values for the new time step are computed from the previous time steps according to App. B.
@@
Figure 5: Imaginary time evolution of SU(2) gauge theory on a periodic 12-plaquette ladder. The Hamiltonian is Eq. (9) with gauge coupling $g=1$. The true ground state is obtained successfully because of the use of self-mitigation. The code ran on 12 qubits of ibm_brisbane.
@@
In the quantum computation of Fig. 5, each quantum circuit used between 40 and 50 randomized compilings with 1000 shots per compiling, and error bars represent 95% confidence intervals. The time step is $\Delta\tau=0.1$. Without self-mitigation, only the first step moves toward the true result and then all subsequent steps move in the wrong direction, heading toward pure noise. With self-mitigation, early time steps move toward the true result and later time steps remain there, exactly as they should when hardware errors are mitigated successfully.
@@ V Real-time evolution
which represents a bare torelon at rest. Evolution of $\left|J_{0}\right>$ through real time will show that it is a superposition of several eigenstates. The dominant eigenstates correspond to the separation of the original diagonal torelon (6 gauge links) into one horizontal and one vertical torelon (4 gauge links each). The dynamics can be visualized from Fig. 2(d) and numerical results are given in Fig. 6. The calculation was done by using the exact eigenvalues and eigenvectors determined in Sec. III. Note that all data in Fig. 6 are translation invariant even through the legend shows only one term for brevity.
Figure 6: Real-time evolution of SU(2) gauge theory on a triamond lattice of three unit cells (12 qubits) and gauge coupling given by $g^{-4}=2$. The initial state is a bare diagonal torelon at rest. Time evolution shows the probabilities of separation into two separate torelons, one horizontal and one vertical.
@@
Consider now the possibility of shrinking the lattice of Fig. 2 from its length of 3 unit cells down to a length of 2.75 unit cells. Specifically, keep the first 11 square paths of Fig. 2(b) and erase the 12th one. The periodic boundary condition on the flattened map now represents a twisted boundary condition for the triamond lattice. For example, a horizontal torelon that travels completely around the long $\hat{j}$ direction of the lattice will return to its original position as a vertical torelon rather than remaining horizontal.
@@
Figure 7: False vacuum decay on a triamond lattice. The lattice has $2.75$ unit cells (11 qubits) and an SU(2) gauge coupling given by $g^{-4}=2$. The initial state is a bare diagonal torelon at rest. Real-time evolution shows the probabilities of separation into two separate torelons, and it also shows the probability of decay into the true vacuum state with no torelons remaining.
The initial state of one diagonal torelon with zero momentum is our false vacuum. It would be completely stable except for the lattice’s twisted boundary condition. The true vacuum is the state containing no torelons at all. To make the transition from the false vacuum to the true vacuum, the original torelon needs to separate into a pair (which is a higher energy state) and the two parts must traverse the entire lattice by travelling in opposite directions. The growing distance between the two parts is clear from the ordering of the curves in the lower left corner of Fig. 7. The significant energy barrier between the false vacuum and the true vacuum results in a transition probability of about 5% according to Fig. 7.
@@ Acknowledgements.
Appendix A Energy eigenvalues and eigenvectors
@@
Table 1: The smallest energy eigenvalues in the no-torelon sector with their eigenvectors, for $g=1$. Only the leading contributions to each eigenvector are listed. Entries containing $\pm$ are degenerate pairs. The eigenvalues are displayed graphically in Fig. 3.
@@
$7.93$ $0.23\left|B_{\pi}\right>-0.23\left|E_{\pi}\right>+\ldots$
$7.89$ $0.16\left|A\right>-0.23\left|B_{0}\right>+0.23\left|E_{0}\right>+\ldots$
$7.69$ $0.80\left|B_{\pm 2\pi/3}\right>-0.56\left|E_{\pm 2\pi/3}\right>+\ldots$
$7.66$ $0.46\left|B_{\pm\pi/2}\right>-0.86\left|D_{\pm\pi/2}\right>+\ldots$
$7.64$ $0.86\left|B_{\pm\pi/2}\right>+0.46\left|D_{\pm\pi/2}\right>+\ldots$
$7.60$ $0.78\left|B_{\pm\pi/3}\right>-0.57\left|D_{\pm\pi/3}\right>+\ldots$
$6.95$ $0.39\left|B_{\pm 2\pi/3}\right>-0.70\left|D_{\pm 2\pi/3}\right>-0.57\left|E_{\pm 2\pi/3}\right>+\ldots$
$6.94$ $0.70\left|B_{\pm 5\pi/6}\right>+0.69\left|D_{\pm 5\pi/6}\right>+\ldots$
$6.91$ $0.45\left|B_{\pm\pi/3}\right>+0.69\left|D_{\pm\pi/3}\right>-0.53\left|E_{\pm\pi/3}\right>+\ldots$
$6.86$ $0.72\left|B_{\pm\pi/6}\right>+0.67\left|D_{\pm\pi/6}\right>+\ldots$
$5.90$ $0.41\left|B_{\pi}\right>+0.69\left|D_{\pi}\right>+0.79\left|E_{\pi}\right>+\ldots$
@@
$-0.67$ $0.96\left|A\right>+0.21\left|B_{0}\right>-0.16\left|C_{0}\right>+\ldots$
@@
[…]
Table 2: The smallest energy eigenvalues in the horizontal torelon sector with their eigenvectors, for $g=1$. Only the leading contributions to each eigenvector are listed. Entries containing $\pm$ are degenerate pairs. The eigenvalues are displayed graphically in Fig. 3.
value vector
$7.70$ $0.97\left|G_{\pi}\right>+\ldots$
$7.66$ $0.97\left|G_{\pm 2\pi/3}\right>+\ldots$
$7.60$ $0.94\left|G_{\pm\pi/3}\right>+\ldots$
[…]
$4.49$ $0.97\left|F_{\pi}\right>+\ldots$
$4.12$ $0.98\left|F_{\pm 2\pi/3}\right>+\ldots$
$3.11$ $0.98\left|F_{\pm\pi/3}\right>+\ldots$
$2.42$ $0.96\left|F_{0}\right>+0.12\left|G_{0}\right>+\ldots$
@@
Table 3: The smallest energy eigenvalues in the diagonal torelon sector with their eigenvectors, for $g=1$. Only the leading contributions to each eigenvector are listed. Entries containing $\pm$ are degenerate pairs. The eigenvalues are displayed graphically in Fig. 3.
@@
$6.13$ $0.31\left|J_{0}\right>+0.45\left|K_{0}\right>+0.81\left|L_{0}\right>+\ldots$
$6.11$ $0.97\left|J_{\pm\pi/2}\right>+\ldots$
$5.99$ $0.97\left|J_{\pm 2\pi/3}\right>+\ldots$
$5.73$ $0.96\left|J_{\pm\pi/3}\right>+\ldots$
$5.56$ $0.95\left|J_{\pm 5\pi/6}\right>+\ldots$
$5.34$ $0.96\left|J_{\pi}\right>+\ldots$
$5.06$ $0.95\left|J_{\pm\pi/6}\right>+\ldots$
$4.66$ $0.92\left|J_{0}\right>+\ldots$
Appendix B Coefficients for the QITE algorithm
@@
The computations in this appendix have been implemented successfully on ibm_brisbane as discussed in Sec. IV.

### v2

@@ I Introduction
The classic picture of false vacuum decay is a double-well potential energy function where the shallower well (called the false vacuum) is separated from the deeper well (called the true vacuum) by an energy barrier [1, 2]. A system that begins in the false vacuum can later be found in the true vacuum with some probability. For the corresponding scenario in a quantum field theory, the decay rate can be calculated by using a semiclassical approximation and imaginary time, but that approach does not explain the physical mechanism of the decay.
In contrast, real-time simulations would allow direct observation of the vacuum decay mechanism, perhaps showing bubbles of true vacuum appearing and nucleating to displace the false vacuum. Various approaches are being explored by several authors [3, 4, 5, 6, 7, 8], including quantum simulations of Ising models [9, 10, 11, 12, 13] and the Schwinger model [14]. In the present work, we take a first step toward using a non-Abelian lattice gauge theory for the real-time simulation of false vacuum decay. We use a pure gauge theory; no matter fields are required.
@@
A promising approach to real-time physics is to rewrite the theory as a Hamiltonian problem, keeping time as a continuous variable and discretizing only the spatial components onto a three-dimensional lattice [17]. The length of the state vector grows exponentially with the lattice size but, in principle, it can be stored efficiently in the qubit register of a quantum computer [18]. Unfortunately, the non-Abelian version of Gauss’s law must now be imposed as a constraint that is external to the Hamiltonian, and finding the best approach to this issue is a high-priority topic for several research groups [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Nevertheless, non-Abelian theories have already been put successfully onto quantum computers [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], though the lattices are necessarily small and typically have just one or two spatial dimensions.
@@
In the absence of matter fields, gauge invariance requires any flow of non-Abelian charges (such as the colors of QCD) to form closed paths on the lattice. Localized particles called glueballs are gauge-invariant objects described by a superposition of localized paths. If the lattice has periodic boundary conditions, there can also be paths that go all the way around the lattice to close on themselves from behind as shown in Fig. 1, and these are called torelons [57, 58, 59, 60, 61, 62, 63, 64]. A torelon is a closed path of color flux that is not homotopic to a point. Torelons become nondynamical heavy states as the lattice volume is taken to infinity, but on small lattices they are central to the physics.
Figure 1: Closed path of round-the-world color flux, called a torelon, in a cubic spatial volume with periodic boundary conditions. (a) One horizontal torelon. (b) One vertical torelon. (c) One diagonal torelon.
@@ II The triamond lattice
A primary motivation for choosing a triamond lattice rather than a cubic lattice is to reduce the number of gauge links touching each site from 6 to 3 [46]. Having only three gauge links at a site means the combined SU(2) value of any pair of links is determined uniquely by the value of the third gauge link. In contrast, the combined SU(2) value of two links on a cubic lattice is not unique because the other four links can be combined in many ways, and this means extra qubits would be required at each lattice site to fully specify the quantum state.
@@
Figure 2: (a) Triamond lattice comprising three unit cells with periodic boundary conditions in all three directions. Twelve of the 36 gauge links are numbered so they can be referenced in the text. (b) A flattened map of the same triamond lattice with the same 12 labels. (c) The same flattened map with two plaquettes shown in thick lines, namely a 10-sided plaquette at 1,2,3,4 and an 8-sided plaquette at 9,11. (d) Thick black paths show a horizontal, diagonal and vertical torelon in the left, center and right unit cells, respectively.
@@
On our small triamond lattice, some of the elementary plaquettes wrap around a periodic boundary and affect the same link twice. In the special case where the twice-affected link retains its original value, the result is pair creation of two torelons that we will refer to as an 8-sided plaquette. Figure 2(c) shows examples of 10-sided and 8-sided plaquettes.
@@
The mapping of a triamond lattice to a ladder of square paths is quite convenient. The three-dimensional aspects of the triamond lattice appear through the 10-sided and 8-sided plaquettes that involve more than just nearest-neighbor interactions among the squares.
@@ III The torelon spectrum
Figure 3: Smallest eigenvalues of SU(2) gauge theory on a triamond lattice having three unit cells in a row and periodic boundary conditions. The gauge theory is truncated such that each gauge link has $j\in\{0,\frac{1}{2}\}$ and the gauge coupling is $g=1.3$. Numerical values from this graph are provided in Appendix A.
@@
The 12 lowest energy states in the diagonal torelon sector tell an important story about how momentum is handled on a triamond lattice. These 12 lowest states are the energy eigenstates that correspond to placing a diagonal torelon at any of the 12 allowed positions on the lattice. In the ladder map of Fig. 2(b), they appear as superpositions of a single six-link rectangle at each of the 12 possible locations. Notice that half of these rectangles are doubly blue and nonyellow while the other half are nonblue and doubly yellow.
The translation along the 12-step ladder of Fig. 2(b) involves a $\pi/2$ rotation with each step along the $\hat{j}$ direction and is called a screw translation. The Fourier transform of this screw translation results in eigenstates of screw momentum that are also eigenstates of energy, and these are precisely the states shown in Fig. 3. Importantly, the 12 values of screw momentum contain six values of linear momentum for the nonyellow torelons and six values for the nonblue torelons. On the original triamond lattice, the nonblue torelon lies along a lattice diagonal that is spatially orthogonal to the nonyellow torelon. A nonblue torelon path is sketched in Fig. 2(d).
@@ IV Imaginary time evolution
Although the triamond lattice for three unit cells can be mapped onto the ladder of Fig. 2(b), the three-dimensional nature of the triamond structure remains evident through the large plaquettes shown in Fig. 2(c). A simpler system is obtained by using the ladder directly as a quasi-one-dimensional lattice having only square plaquettes and no connection to the three-dimensional triamond structure. This ladder of simple square plaquettes has been used in several quantum computations for SU($N$) gauge theory [33, 34, 35, 36, 37]. Imaginary time evolution [68] for a nonperiodic ladder of only three rungs was run on a quantum computer recently [46]. Here we will study the 12-rung periodic ladder.
@@
For the 12-qubit Hamiltonian of Eq. (9), the most general expression has many terms but we might expect only a subset of them to dominate the physics. In particular, since $H_{\square}$ has only single-qubit terms, adjacent-pair terms, and adjacent-triple terms, we anticipate that $A$ will be predominantly local. Specifically, the single-qubit and adjacent-pair terms are expected to be most important, with adjacent-triple terms being less important. Since adjacent-triple terms would require swap gates when implemented on IBM’s heavy-hex hardware architecture, we propose to neglect these subleading terms. Therefore our ansatz has only single-qubit and adjacent-pair terms,
@@
Notice that each qubit $j$ gets its own set of coefficients $(a_{\bullet})_{j}$, thus allowing our error mitigation to make no assumption about individual physical qubits having similar noise profiles. The coefficients $(a_{\bullet})_{j}$ are determined from state tomography as explained in detail in Appendix B. The exponentiation of each term in $A$ is accomplished by using
@@
and their numerical values are different at each time step in the quantum circuit. Values from previous time steps are stored in a list for reuse, and values for the new time step are computed from the previous time steps according to Appendix B.
@@
Figure 5: Imaginary time evolution of SU(2) gauge theory on a periodic 12-plaquette ladder. The Hamiltonian is Eq. (9) with gauge coupling $g=1$. The true ground state is obtained successfully because of the use of self-mitigation. The code ran on 12 qubits of ibm_brisbane. Numerical values from this graph are provided in Table 6.
@@
In the quantum computation of Fig. 5, each quantum circuit used between 40 and 50 randomized compilings with 1000 shots per compiling, and error bars represent 95% confidence intervals. The time step is $\Delta\tau=0.1$. Without self-mitigation, only the first step moves toward the true result and then all subsequent steps move in the wrong direction, heading toward pure noise. With self-mitigation, early time steps move toward the true result and later time steps remain there, exactly as they should when hardware errors are mitigated successfully. Code that ran on ibm_brisbane is publicly available [82]. Each time step used about 100 seconds of time on the quantum hardware for the combined running of the physics and self-mitigation circuits with all randomized compilings.
@@ V Real-time evolution
which represents a bare torelon at rest. Evolution of $\left|J_{0}\right>$ through real time will show that it is a superposition of several eigenstates. The dominant eigenstates correspond to the separation of the original diagonal torelon (six gauge links) into one horizontal and one vertical torelon (four gauge links each). The dynamics can be visualized from Fig. 2(d) and numerical results are given in Fig. 6. The calculation was done by using the exact eigenvalues and eigenvectors determined in Sec. III. Note that all data in Fig. 6 are translation invariant even through the legend shows only one term for brevity.
Figure 6: Real-time evolution of SU(2) gauge theory on a triamond lattice of three unit cells (12 qubits) and gauge coupling $g=1.3$. The initial state is a bare diagonal torelon at rest. Time evolution shows the probabilities of separation into two separate torelons, one horizontal and one vertical. Numerical values from this graph are provided in Table 4.
@@
Consider now the possibility of shrinking the lattice of Fig. 2 from its length of three unit cells down to a length of 2.75 unit cells. Specifically, keep the first 11 square paths of Fig. 2(b) and erase the 12th one. The periodic boundary condition on the flattened map now represents a twisted boundary condition for the triamond lattice. For example, a horizontal torelon that travels completely around the long $\hat{j}$ direction of the lattice will return to its original position as a vertical torelon rather than remaining horizontal.
@@
Figure 7: False vacuum decay on a triamond lattice. The lattice has $2.75$ unit cells (11 qubits) and an SU(2) gauge coupling $g=1.3$. The initial state is a bare diagonal torelon at rest. Real-time evolution shows the probabilities of separation into two separate torelons, and it also shows the probability of decay into the true vacuum state with no torelons remaining. Numerical values from this graph are provided in Table 5.
The initial state of one diagonal torelon with zero momentum is our false vacuum. It would be completely stable except for the lattice’s twisted boundary condition. The true vacuum is the state containing no torelons at all. To make the transition from the false vacuum to the true vacuum, the original torelon needs to separate into a pair (which is a higher energy state) and the two parts must traverse the entire lattice by travelling in opposite directions. The growing distance between the two parts is clear from the ordering of the curves in the lower left corner of Fig. 7. The significant energy barrier between the false vacuum and the true vacuum results in a transition probability of about 2% according to Fig. 7.
@@ Acknowledgements.
Data Availability
The data supporting this study’s findings are available within the article.
APPENDIX A ENERGY EIGENVALUES
AND EIGENVECTORS
@@
Table 1: The smallest energy eigenvalues in the no-torelon sector with their eigenvectors, for $g=1.3$. Only the leading contributions to each eigenvector are listed. Entries containing $\pm$ are degenerate pairs. The eigenvalues are displayed graphically in Fig. 3.
@@
7.9907 $0.22\left|A\right>-0.76\left|B_{0}\right>+0.57\left|E_{0}\right>+\ldots$
7.8551 $0.23\left|B_{\pi}\right>-0.57\left|E_{\pi}\right>+\ldots$
7.5866 $0.79\left|B_{\pm 2\pi/3}\right>+0.56\left|E_{\pm 2\pi/3}\right>+\ldots$
7.5673 $0.15\left|B_{\pm\pi/2}\right>+0.96\left|D_{\pm\pi/2}\right>+\ldots$
7.5224 $0.96\left|B_{\pm\pi/2}\right>-0.15\left|D_{\pm\pi/2}\right>+\ldots$
7.4997 $0.78\left|B_{\pm\pi/3}\right>-0.57\left|D_{\pm\pi/3}\right>+\ldots$
6.5693 $0.40\left|B_{\pm 2\pi/3}\right>-0.70\left|D_{\pm 2\pi/3}\right>-0.56\left|E_{\pm 2\pi/3}\right>+\ldots$
6.5359 $0.44\left|B_{\pm\pi/3}\right>-0.69\left|D_{\pm\pi/3}\right>-0.54\left|E_{\pm\pi/3}\right>+\ldots$
6.5357 $0.70\left|B_{\pm\pi/6}\right>-0.69\left|D_{\pm\pi/6}\right>+\ldots$
6.5133 $0.70\left|B_{\pm 5\pi/6}\right>-0.68\left|D_{\pm 5\pi/6}\right>+\ldots$
@@
5.1084 $0.42\left|B_{\pi}\right>+0.68\left|D_{\pi}\right>+0.55\left|E_{\pi}\right>+\ldots$
[…]
@@
$\left|H_{1}\right>$ $=$ $\frac{1}{\sqrt{2}}(\left|110010000000\right>+\left|100110000000\right>),~{}~{}~{}~{}~{}~{}$ (50)
$\left|I_{1}\right>$ $=$ $\frac{1}{\sqrt{2}}(\left|110000100000\right>+\left|100001100000\right>),~{}~{}~{}~{}~{}~{}$ (51)
Table 2: The smallest energy eigenvalues in the horizontal torelon sector with their eigenvectors, for $g=1.3$. Only the leading contributions to each eigenvector are listed. Entries containing $\pm$ are degenerate pairs. The eigenvalues are displayed graphically in Fig. 3.
value vector
7.5222 $0.96\left|G_{\pi}\right>$
7.4834 $0.94\left|G_{\pm 2\pi/3}\right>+\ldots$
7.3943 $0.88\left|G_{\pm\pi/3}\right>-0.46\left|H_{\pm\pi/3}\right>$
@@
4.7162 $0.96\left|F_{\pi}\right>+\ldots$
4.1822 $0.97\left|F_{\pm 2\pi/3}\right>+\ldots$
2.8189 $0.97\left|F_{\pm\pi/3}\right>+\ldots$
1.8586 $0.95\left|F_{0}\right>+\ldots$
@@
Table 3: The smallest energy eigenvalues in the diagonal torelon sector with their eigenvectors, for $g=1.3$. Only the leading contributions to each eigenvector are listed. Entries containing $\pm$ are degenerate pairs. The eigenvalues are displayed graphically in Fig. 3.
@@
7.8411 $0.58\left|K_{\pm\pi/2}\right>+0.80\left|L_{\pm\pi/2}\right>+\ldots$
7.5380 $0.79\left|K_{\pm\pi/2}\right>-0.57\left|L_{\pm\pi/2}\right>+\ldots$
7.1800 $0.81\left|K_{\pm 2\pi/3}\right>-0.51\left|L_{\pm 2\pi/3}\right>+\ldots$
7.1496 $0.06\left|J_{\pm\pi/6}\right>+0.79\left|K_{\pm\pi/6}\right>+0.51\left|L_{\pm\pi/6}\right>+\ldots$
7.1183 $0.79\left|K_{\pi}\right>+0.50\left|L_{\pi}\right>+\ldots$
6.6515 $0.48\left|K_{\pm\pi/3}\right>+0.84\left|L_{\pm\pi/3}\right>+\ldots$
5.9206 $0.63\left|J_{\pm 5\pi/6}\right>-0.31\left|K_{\pm 5\pi/6}\right>-0.67\left|L_{\pm 5\pi/6}\right>+\ldots$
5.8511 $0.97\left|J_{\pm\pi/2}\right>+\ldots$
5.8075 $0.96\left|J_{\pm 2\pi/3}\right>+\ldots$
5.6195 $0.51\left|J_{0}\right>+0.37\left|K_{0}\right>+0.75\left|L_{0}\right>+\ldots$
5.4622 $0.95\left|J_{\pm\pi/3}\right>+\ldots$
5.3389 $0.92\left|J_{\pi}\right>+\ldots$
5.3357 $0.72\left|J_{\pm 5\pi/6}\right>+0.43\left|K_{\pm 5\pi/6}\right>-0.49\left|L_{\pm 5\pi/6}\right>+\ldots$
4.8470 $0.93\left|J_{\pm\pi/6}\right>-0.27\left|K_{\pm\pi/6}\right>+\ldots$
[…]
Evolution through real time can be calculated from the full set of eigenvalues and eigenvectors. For 3 unit cells, results are shown numerically in Table 4 and graphically in Fig. 6. For 2.75 unit cells, results are shown numerically in Table 5 and graphically in Fig. 7.
Table 4: Time evolution of a diagonal torelon at rest on 3 unit cells. Labels $J$, $K$ and $L$ correspond to Eqs. (52-57). Results from this table are displayed graphically in Fig. 6.
time probability
$J$ $K$ $L$
0.25 0.8067 0.0304 0.0028
0.50 0.7343 0.0824 0.0356
0.75 0.7025 0.1309 0.1117
1.00 0.5860 0.0902 0.2035
1.25 0.5334 0.0540 0.3011
1.50 0.4169 0.1070 0.3281
1.75 0.3271 0.2225 0.3208
2.00 0.2256 0.3620 0.3273
2.25 0.1658 0.3443 0.3468
2.50 0.1649 0.2645 0.4356
2.75 0.1965 0.1569 0.5142
3.00 0.2341 0.1246 0.5041
3.25 0.3090 0.2103 0.4171
3.50 0.3555 0.2344 0.2720
3.75 0.4451 0.2291 0.1511
4.00 0.6814 0.1395 0.1187
Table 5: Time evolution of a diagonal torelon at rest on 2.75 unit cells. Column headings 10, 8, 6, 4, 2, 0 correspond to the steps remaining until torelon annihilation. Results from this table are displayed graphically in Fig. 7.
time probability
10 8 6 4 2 0
0.25 0.8276 0.0320 0.0024 0.0001 0.0000 0.0000
0.50 0.7470 0.0976 0.0227 0.0036 0.0005 0.0000
0.75 0.7232 0.1550 0.0650 0.0132 0.0011 0.0002
1.00 0.6028 0.1439 0.1057 0.0356 0.0048 0.0008
1.25 0.5564 0.1133 0.1190 0.0767 0.0267 0.0032
1.50 0.5280 0.0965 0.0941 0.1139 0.0533 0.0073
1.75 0.5058 0.0999 0.0717 0.1158 0.0895 0.0131
2.00 0.4685 0.0973 0.0720 0.1122 0.1198 0.0173
2.25 0.4494 0.0790 0.0867 0.1249 0.1226 0.0166
2.50 0.4268 0.0765 0.1269 0.1639 0.0953 0.0128
2.75 0.3823 0.0974 0.1642 0.1812 0.0503 0.0082
3.00 0.3453 0.1516 0.2019 0.1593 0.0140 0.0048
3.25 0.3222 0.2279 0.2396 0.1156 0.0108 0.0038
3.50 0.3215 0.2720 0.2184 0.0595 0.0106 0.0028
3.75 0.3581 0.2667 0.1751 0.0361 0.0126 0.0026
4.00 0.4960 0.2191 0.1474 0.0634 0.0122 0.0027
APPENDIX B COEFFICIENTS FOR THE
QITE ALGORITHM
@@
The computations in this appendix have been implemented successfully on ibm_brisbane as discussed in Sec. IV. Numerical results are provided in Fig. 5 and Table 6. Our quantum computer code for the QITE algorithm is publically available [82].
Table 6: Imaginary time evolution on 12 qubits of ibm_brisbane. The Hamiltonian is Eq. (9) with gauge coupling $g=1$ on a periodic 12-plaquette ladder. The true ground state is obtained successfully through self-mitigation. The eigenvalues are displayed graphically in Fig. 5.
$\tau$ $\psi\right|H_{\square}\left|\psi\right>$
without self-mitigation with self-mitigation
0.1 -3.15$\pm$0.19 -6.51$\pm$0.20
0.2 -2.00$\pm$0.19 -8.46$\pm$0.28
0.3 -0.65$\pm$0.20 -8.92$\pm$0.35
0.4 1.09$\pm$0.23 -9.65$\pm$0.37
0.5 2.79$\pm$0.22 -9.08$\pm$0.50
0.6 4.20$\pm$0.21 -9.38$\pm$0.58
0.7 5.15$\pm$0.15 -9.57$\pm$0.47
0.8 7.32$\pm$0.17 -8.59$\pm$0.62
0.9 7.85$\pm$0.19 -9.22$\pm$0.85